Translate the following into a system of equations, then solve it: A customer walks into an electronics store and buys 5 MP3 players and 8 sets of headphones, paying $840. A 2nd customer buys 3 MP3 players and 4 sets of headphones, and pays $480. How much does an MP3 player cost? How much does a set of headphones cost?: Translate the following into a system of equations, then solve it: A customer walks into an electronics store and buys 5 MP3 players and 8 sets of headphones, paying $840. A 2nd customer buys 3 MP3 players and 4 sets of headphones, and pays $480. How much does an MP3 player cost? How much does a set of headphones cost?
Answer by checkley77(3397)

5MP+8H=840
3MP+4H=480 MULTIPLY THIS EQUATION BY -2 & ADD THEM
-6MP-8H=-960
5MP+8H=840
----------------------
-MP=-120
MP=$120 COST OF EACH MP3 PLAYER.
5*120+8H=840
600+8H=840
8H=840-600
8H=240
H=240/8
H=$30 COST OF EACH SET OF HEADPHONES.
PROOF:
3*120+4*30=480
360+120=480
480=480